OpenStudy (mlgpro):

Casey and Jeremiah made a simple bike ramp out of a piece of straight plywood that is 36 inches long. If the top of the ramp stands 18 inches high, what is the length of the ramp along the ground? Round your answer to the nearest inch. 25 in 31 in 36 in 40 in

OpenStudy (sedatefrog712):

X^2 + 18^2 = 36^2 X^2 + 324 = 1296 X^2 = 972 X = 31

OpenStudy (mokeira):

 Use Pythagoras